I am totally lost on this two question and need some help.
Solve and check the following eqaution for x:
x/x-5+2=5/x-5
and
5/x+6+2/x^2+7x+6=3/x+1 Thank you

Answer by stanbon(75887)

Solve and check the following eqaution for x:
x/(x-5) +2 = 5/(x-5)
Multiply both sides by (x-5) to get:
x + 2(x-5) = 5
x + 2x -10 = 5
3x= 15
x=5
-------
But x cannot be 5 because that would make the denominator of the
original equation zero.
-----------
Final Answer: No solutions
==================

5/(x+6) + 2/(x^2+7x+6) = 3/(x+1)
Multiply thru by (x+6)(x+1) which is the same as x^2+7x+6
---------
5(x+1) + 2 = 3(x+6)
5x+5+2 = 3x+18
5x+7 = 3x+18
2x = 11
x= 11/2
--------
Cheers,
Stan H.